# Lensless fiber endomicroscopic phase imaging with speckle-conditioned diffusion model

### v1

@@ 2.5 Endomicroscopic phase reconstruction of tissue samples
Based on the model pretrained on ImageNet, we finetune it on $150$ tissue images, and the reconstructed images of U-Net and SpecDiffusion are depicted in Fig. 6c and d. SpecDiffusion clearly outlines the cell morphology and boundaries, facilitating precise cell analysis and enumeration which are crucial for b. Moreover, it vividly captures cellular sociology, depicting the cellular aggregation patterns within tissues. The MAE, PSNR, SSIM and 2D correlation coefficient distributions of U-Net and SpecDiffusion are illustrated in Fig. 6e-h. The distributions of SpecDiffusion are concentrated in regions indicative of better metric performance and exhibit a more centered tendency on most metrics. This suggests that the phase reconstruction capacity of SpecDiffusion is reliable and consistent. The averaged evaluation metrics are summarized in Tab. 1. Quantitatively, SpecDiffusion reduces MAE from $0.2091~{}rad$ to $0.0304~{}rad$, indicating its accurate reconstruction. It also demonstrates SpecDiffusion’s potential for further applications in measuring biophysical properties, such as refractive index and dry mass. Meanwhile, PSNR is improved from $19.43~{}dB$ to $34.98~{}dB$, suggesting that SpecDiffusion extracts more information from speckles. SpecDiffusion also achieves higher SSIM from $0.4310$ to $0.9697$, which is an indicator for retaining more structural information. The elevated 2D correlation coefficient from $0.6999$ to $0.9915$ demonstrates a greater overall fidelity between the reconstructed images and their ground truth. These metrics collectively affirm that, SpecDiffusion’s reconstructions are high-fidelity and reliable, thereby demonstrating its applicability to complex image reconstruction tasks. More tissue reconstruction results and the corresponding residual maps are illustrated in supplementary Fig. S3333. The residual maps demonstrate low reconstruction errors for SpecDiffusion, verifying its phase reconstruction precision. This enhanced performance and precision makes SpecDiffusion a promising tool for advanced medical imaging and diagnostics, offering significant benefits for clinical practices.
@@
Figure 7: Cell segmentation by SAM on ground truth and reconstructed images. a Cell segmentation based on lensless MCF phase imaging by SAM. b Ground truth phase images and their cell segmentation results. c and d Reconstructed phase images by U-Net and SpecDiffusion, and their cell segmentation results. Scale bars $50~{}\mu m$
@@
To further assess the potential of our reconstructed images in conjunction with other AI technologies, we employ the segmentation anything model (SAM) [44] for zero-shot cell segmentation tasks on reconstructed and ground truth images, as illustrated in Fig. 7a. SAM stands out as an AI segmentation model characterized by its exceptional generalization ability, capable of delivering outstanding segmentation outcomes without the prior knowledge about the samples. The details for segmentation task is explained in Materials and Methods section. As AI technologies like SAM are increasingly integrated into medical diagnostics, it’s important to ensure that the reconstructed images maintain fidelity to their originals within the realm of modern AI technology.
The segmentation results of SpecDiffusion’s reconstructed images are illustrated in Fig. 7d, which closely mirror those of the ground truth in Fig. 7b, accurately depicting cell locations, sizes and morphologies. This performance starkly contrasts with U-Net, whose segmentation results, indicated in Fig. 7c, are limited to identifying the general boundaries of different tissues without providing the granularity required for detailed cell-level segmentation. To quantitatively evaluate the segmentation quality, we employ the intersection over union (IoU) to measure the agreement between segmentation results derived from reconstructed phase images and their ground truth counterparts. The methodology for calculating IoU is detailed in Materials and Methods section. The averaged IoU is illustrated in Tab. 1. SpecDiffusion substantially enhances segmentation accuracy, improving the IoU from $0.2480$ to $0.7658$. The distribution of IoU values, highlighted in supplementary Fig. S5555, indicates the aggregation of SpecDiffusion’s segmentation results within the higher IoU range, suggesting that the majority of its reconstructions are conducive to effective segmentation. The results signify the alignment between SpecDiffusion’s reconstructed images and the ground truth in the context of segmentation tasks, indicating its higher fidelity for downstream applications.
@@
Method MAE(rad) PSNR SSIM 2D Correlation IoU
U-Net 0.2091 19.43 0.4310 0.6999 0.2480
SpecDiffusion 0.0304 34.98 0.9697 0.9915 0.7658
3 Discussion
We demonstrate a novel conditional diffusion model SpecDiffusion, which use the intensity-only speckle images to guide the diffusion model in each step, thereby facilitating high-resolution phase reconstruction of complex objects through lensless fiber endomicroscope. As the emerging model in AI-driven image generation, diffusion models are skilled at reconstructing complex and realistic images with meticulous details, providing a powerful tool for enhancing optical imaging capabilities. With the support of diffusion mechanism, SpecDiffusion achieves higher resolution capability and maintains details better for complex images than conventional methods, such as U-Net. This enables high resolution imaging in tiny regions using the lensless endomicroscope, crucial for medical analysis and disease diagnosis. Compared to typical diffusion models, SpecDiffusion leverages speckle as supervisor and is trained from scratch in a supervised way, effectively improving the reconstruction fidelity and paving the path to integrate diffusion models into optics imaging. Furthermore, our method exhibits potential for specialized practical applications. By utilizing a minimal set of tissue images for transfer learning, SpecDiffusion rapidly gets itself adapted and delivers clear and detailed phase reconstructions, indicating its powerful generalization capacity in various application scenarios. Additionally, when applied to segmentation tasks using SAM, images reconstructed by SpecDiffusion yields results that closely align with those of ground truth, revealing the feasibility of SpecDiffusion-reconstructed images in down-stream applications.
Unlike traditional digital holography for quantitative phase imaging [5], our speckle reconstruction method offers simplicity and flexibility, facilitating real-time cell observation without calibration. This allows biological researchers to swiftly gain insight into their samples without the need for complex preparations. Thus, we provide a promising tool to quickly verify the effectiveness of their medicine, or observe the biochemical reactions inside the sample.
The iterative process of SpecDiffusion might raises concern about its time efficiency. Executed on Pytorch with an NVIDIA A100100100100 GPU, the reconstruction process achieves a rate of about $1.77$ FPS. In comparison, the traditional iterative method FAST [5] requires roughly $24$ seconds per image after calibration. In clinical settings, the reconstruction speed of SpecDiffusion speed is adequate for doctors to assess tissue effectively. Furthermore, the utilization of parallel computation across multiple GPUs can significant accelerate the reconstruction process, providing faster results when needed.
4 Materials and Methods
4.1 SpecDiffusion model
@@ 4.2 Evaluation metrics
$\text{MAE}(I,I^{\prime})=\frac{1}{mn}\sum\limits_{i=0}^{m-1}\sum\limits_{j=0}^{n-1}|I(i,j)-I^{\prime}(i,j)|.$ (3)
@@
$\text{MSE}(I,I^{\prime})=\frac{1}{mn}\sum\limits_{i=0}^{m-1}\sum\limits_{j=0}^{n-1}[I(i,j)-I^{\prime}(i,j)]^{2}.$ (5)
@@
[…]
Funding This work is supported by the Shanghai AI Laboratory, National Key R&D Program of China (2022ZD0160100), the National Natural Science Foundation of China (62376222), and Young Elite Scientists Sponsorship Program by CAST (2023QNRC001).
[…]
Acknowledgments The authors would like to thank the valuable support from Prof. Juergen Czarske and BIOLAS in TU Dresden.
[…]
Disclosures The authors declare no conflicts of interest.
[…]
Data Availability Statement SpecDiffusion model and the test codes are available at  [46, 47, 48].
[…]
Supplemental document See Supplement 1 for supporting content.

### v4

@@ Results
Based on the model pretrained on ImageNet, we finetune it on $150$ tissue images, and the reconstructed images of U-Net and SpecDiffusion are depicted in Fig. 5c and d. SpecDiffusion clearly outlines the cell morphology and boundaries, facilitating precise cell analysis and enumeration which are crucial for b. Moreover, it vividly captures cellular sociology, depicting the cellular aggregation patterns within tissues. The MAE, PSNR, SSIM and 2D correlation coefficient distributions of U-Net and SpecDiffusion are illustrated in Fig. 5e-h. The distributions of SpecDiffusion are concentrated in regions indicative of better metric performance and exhibit a more centered tendency on most metrics. This suggests that the phase reconstruction capacity of SpecDiffusion is reliable and consistent. The averaged evaluation metrics are summarized in Table 1. Quantitatively, SpecDiffusion reduces MAE from $0.2091\leavevmode\nobreak\ rad$ to $0.0304\leavevmode\nobreak\ rad$, indicating its accurate reconstruction. This reduction also highlights SpecDiffusion’s potential for further applications in measuring biophysical properties, such as refractive index, dry mass and optical transmission matrix [11]. As a result, real-time quantitative measurement during clinical operation becomes possible, enabling precise identification of boundaries between healthy and diseased regions, and minimizing patient harm. Meanwhile, PSNR is improved from $19.43\leavevmode\nobreak\ dB$ to $34.98\leavevmode\nobreak\ dB$, suggesting that SpecDiffusion extracts more information from speckles, thereby supporting the success of subsequent virtual staining. SpecDiffusion also achieves higher SSIM from $0.4310$ to $0.9697$, which is an indicator for retaining more structural information, making the reconstructed tissue structures reliable for clinical cancer diagnosis. Furthermore, the elevated 2D correlation coefficient, from $0.6999$ to $0.9915$, and the improved PRSTD value, from $0.2532$ to $0.0422$, demonstrate greater overall fidelity with less reconstruction bias between different regions, further proving the reliability of the tissue images. These metrics collectively affirm that, SpecDiffusion’s reconstructions are high-fidelity and reliable, thereby demonstrating its applicability to complex image reconstruction tasks. More tissue reconstruction results and the corresponding residual maps are illustrated in supplementary Fig. S4444. The residual maps demonstrate low reconstruction errors for SpecDiffusion, verifying its phase reconstruction precision. This enhanced performance and precision makes SpecDiffusion a promising tool for advanced medical imaging and diagnostics, offering significant benefits for clinical practices.
@@
Figure 6: Cell segmentation by SAM on ground truth and reconstructed images. a Cell segmentation based on lensless MCF phase imaging by SAM. b Ground truth phase images and their cell segmentation results. c and d Reconstructed phase images by U-Net and SpecDiffusion, and their cell segmentation results. Scale bars $50\leavevmode\nobreak\ \mu m$
@@
To further assess the potential of our reconstructed images in conjunction with other AI technologies, we employ the segmentation anything model (SAM) [58] for zero-shot cell segmentation tasks on reconstructed and ground truth images, as illustrated in Fig. 6a. SAM stands out as an AI segmentation model characterized by its exceptional generalization ability, capable of delivering outstanding segmentation outcomes without the prior knowledge about the samples. The details for segmentation task is explained in Methods section. As AI technologies like SAM are increasingly integrated into medical diagnostics, it’s important to ensure that the reconstructed images maintain fidelity to their originals within the realm of modern AI technology.
@@
Method MAE(rad) PSNR SSIM 2D Correlation PRSTD IoU
U-Net 0.2091 19.43 0.4310 0.6999 0.2532 0.2480
SpecDiffusion 0.0304 34.98 0.9697 0.9915 0.0422 0.7658
The segmentation results of SpecDiffusion’s reconstructed images are illustrated in Fig. 6d, which closely mirror those of the ground truth in Fig. 6b, accurately depicting cell locations, sizes and morphologies. This performance starkly contrasts with U-Net, whose segmentation results, indicated in Fig. 6c, are limited to identifying the general boundaries of different tissues without providing the granularity required for detailed cell-level segmentation. To quantitatively evaluate the segmentation quality, we employ the intersection over union (IoU) to measure the agreement between segmentation results derived from reconstructed phase images and their ground truth counterparts. The methodology for calculating IoU is detailed in Methods section. The averaged IoU is illustrated in Table 1. SpecDiffusion substantially enhances segmentation accuracy, improving the IoU from $0.2480$ to $0.7658$. The distribution of IoU values, highlighted in supplementary Fig. S6666, indicates the aggregation of SpecDiffusion’s segmentation results within the higher IoU range, suggesting that the majority of its reconstructions are conducive to effective segmentation. The results signify the alignment between SpecDiffusion’s reconstructed images and the ground truth in the context of segmentation tasks, indicating its higher fidelity for downstream applications.
Discussion
Typical diffusion models [59] rely on a sequence of noise-reduction steps based purely on learned data distributions without explicit physics-based guidance, which limits their effectiveness in applications requiring high precision. SpecDiffusion stands out from conventional diffusion models by incorporating speckle patterns as conditioning inputs, enhancing its capability for detailed and accurate image reconstruction in biomedical imaging. By utilizing physics-informed guidance in its denoising process, SpecDiffusion not only outperforms traditional models like U-Net in performance but also meets the specific demands of fiber endomicroscopic imaging more effectively. Furthermore, we have validated the diffusion-driven fiber endomicroscope’s performance in cell segmentation, this capability can be further extended to other digital pathology applications, such as tumor boundary detection, tissue classification, and morphological analysis, with minimum invasiveness.
In the realm of generative models, SpecDiffusion presents notable advantages over established techniques such as generative adversarial network (GANs) [60] and vision transformers [61]. While GANs are adept at generating visually appealing images, they frequently encounter issues related to training stability and convergence, particularly in complex imaging tasks where the preservation of fine details is critical. Conversely, Vision Transformers, despite their efficacy in processing sequential data, demand substantial computational resources and often struggle to efficiently capture spatial hierarchies that are crucial for high-resolution imaging. SpecDiffusion employs an iterative noise reduction process that is both stable and robust, facilitating the high-resolution reconstruction of phases in a physics-informed manner. This approach not only enhances the model’s ability to handle detailed imaging tasks but also mitigates common problems associated with other generative models, such as training instability and excessive computational load. The integration of speckle patterns as a conditioning input further aids in guiding the denoising process, ensuring that SpecDiffusion maintains a high level of accuracy and detail essential for biomedical applications. Furthermore, the SpecDiffusion offers a resolution comparable to the iterative phase retrieval method  [5], while significantly reducing reconstruction time from 24 seconds per image to just 0.56 seconds. This substantial improvement in efficiency highlights SpecDiffusion’s potential for real-time applications in clinical applications. Moreover, Lensless fiber endomicroscope offers substantial advantages for in vivo imaging, allowing for deep tissue visualization with minimal invasiveness. We have demonstrated that, combined with the high-resolution and accurate phase reconstructions provided by SpecDiffusion, this system could be further employed for real-time, label-free imaging of tissues. When integrated with advanced AI models like SAM, the system enables precise cell segmentation and tumor boundary detection, enhancing the accuracy of cancer diagnosis and tissue classification. The minimally invasive nature of the fiber endomicroscope, along with its potential for high-fidelity imaging, makes this approach particularly suited for real-time diagnostics, surgical guidance, and early detection of diseases in hard-to-reach areas, improving both patient outcomes and digital pathology workflows.
Despite these promising results, there are limitations to the current implementation of SpecDiffusion. While the model is trained on a large dataset of 100,000 images, its performance may still be limited when applied to tissue types or imaging conditions not represented in the dataset. Additionally, although the reconstruction speed is a significant improvement, further optimizations may be necessary for real-time clinical applications. Future work could address these limitations by expanding the dataset to include more diverse tissue types and imaging conditions, as well as optimizing the model for faster, real-time applications. Additionally, exploring the integration of SpecDiffusion with other imaging modalities could provide more comprehensive diagnostic capabilities. Overall, SpecDiffusion represents a significant step forward in quantitative phase reconstruction for lensless fiber endomicroscopy, offering a powerful tool for improving precision in medical diagnostics, particularly in cancer detection and digital pathology with minimal invasiveness.
Methods
Experimental setup
We implement the optical setup illustrated in Fig. S1111 to acquire the training images for the diffusion-driven fiber endomicroscope. In this setup, phase images are projected onto an SLM and imaged by the fiber endomicroscope, with the speckle patterns captured by the detection system. A detailed description of the optical setup is provided in supplementary information. During the experiment, the SLM and detection camera were triggered together to ensure the phase images and speckle patterns were captured at the same time, ensuring accurate alignment in the training dataset. The distance between the far-field image plane for capturing speckles and the MCF facet on the detection side was $0.5$ mm.
Training dataset
Utilizing the optical setup, we construct a tailored dataset for QPI in lensless fiber endomicroscopy. The natural complex images from the ImageNet dataset [55] are converted to grayscale images, and further transformed to phase images. These phase images are resized to a resolution of $980\times 980$ pixels and subsequently zero-padded to $1920\times 1080$ pixels to accommodate the display resolution of the SLM. Computer-generated holograms are then calculated and displayed on the SLM to achieve a precise holographic representation of the ground truth image. Compared to MNIST and Fashion-MNIST datasets in our previous study [40], images from ImageNet are more complex and detailed, thus providing a more robust benchmark for evaluating phase reconstruction capabilities in real-world scenarios. In the experiment, we captured $100,000$ paired phase and speckle images, allocating $90,000$ pairs for training, and the remaining $10,000$ pairs for testing.
SpecDiffusion model
@@ Methods
$\text{MAE}(I,I^{\prime})=\frac{1}{mn}\sum\limits_{i=0}^{m-1}\sum\limits_{j=0}^{n-1}|I(i,j)-I^{\prime}(i,j)|,$ (3)
where $m$ and $n$ indicate the row and column index of the image.
@@
$\text{MSE}(I,I^{\prime})=\frac{1}{mn}\sum\limits_{i=0}^{m-1}\sum\limits_{j=0}^{n-1}[I(i,j)-I^{\prime}(i,j)]^{2},$ (5)
where $m$ and $n$ indicate the row and column index of the image.
@@
The PRSTD is employed to quantify the reconstruction bias in the reconstructed images. Typically, an image’s background occupies a substantial and homogeneous area, making it easier to reconstruct accurately. Thus, data-driven models exhibit a reconstruction bias, favoring the background due to its simplicity. In contrast, the foreground, which contains more informative and complex features, is more challenging to be reconstructed and thus neglected. This bias between background and foreground reconstruction quality is captured by PRSTD. For original image $I$ and reconstructed image $I^{\prime}$, the PRSTD is calculated as
$\text{PRSTD}(I,I^{\prime})=\sigma(|I-I^{\prime}|),$ (8)
where $\sigma$ indicates standard variance.
@@
Acknowledgements
This work is supported by the Shanghai AI Laboratory, National Key R&D Program of China (2022ZD0160102), the National Natural Science Foundation of China (62376222), and Young Elite Scientists Sponsorship Program by CAST (2023QNRC001). The authors would like to thank the valuable support from Dr. Nektarios Koukourakis.
Author contributions statement
Z.C., J.S. and B.Z. conceived the experiments, Z.C. and J.S. conducted the experiments, Z.C. and X.Y. analysed the results, Z.C., J.S. wrote the manuscript, all authors reviewed the manuscript.
Disclosures
The authors declare no conflicts of interest.
Data Availability Statement
The SpecDiffusion model and the source code are publicly available on GitHub [63].
Supplemental document
See Supplement information and supplementary video for supporting content.